OpenStudy (anonymous):

WILL MEDAL AND FAN!! What is the solution to the system? line 1 x.y -1.9 3.-3 line 2 x.y -1.-1 2.5 (___,___) plz and thank you also, I never was goodat these. >.<

OpenStudy (anonymous):

@MSxSammiBooxX

OpenStudy (anonymous):

Ah my account is acting up. Let @ganeshie8 or @mathmate help while I go through this again

OpenStudy (anonymous):

So sorry! I can't see your question it is just blank

OpenStudy (anonymous):

ok. thanks thow. :) I can message it to ya thow

OpenStudy (anonymous):

if u don't mind

jaynator495 (jaynator495):

@Joel_the_boss

JoelTheBoss (joel_the_boss):

@Jhannybean @TuringTest @pooja195 Maybe they can help. :) Im history wiz only :) @Jaynator495

OpenStudy (anonymous):

ok thx. :) im a science nerd math is my enemy lol @Joel_the_boss

pooja195 (pooja195):

hey there welcome to openstudy!

OpenStudy (anonymous):

thx @pooja195 could u help me?

pooja195 (pooja195):

well first off what do u think?

OpenStudy (anonymous):

tbh idk

pooja195 (pooja195):

ok well u know to substitue for 0 everytime u find X and Y right?

OpenStudy (anonymous):

yes

pooja195 (pooja195):

so for -1,9 say u substitue for X what would your Y be?

OpenStudy (anonymous):

Back! I am so so so sorry! Thanks Pooja :P I will let her finish off :) I PMed you back science, did you get it?

OpenStudy (anonymous):

no I haven't @MSxSammiBooxX

OpenStudy (anonymous):

Hmm, OK. Then I will let Pooja continue

pooja195 (pooja195):

its fine :) ok so what would your Y be?

OpenStudy (anonymous):

y would be 3,-3?

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

where did u get that from?

OpenStudy (anonymous):

the line 1

OpenStudy (mathmate):

So you have two lines, L1: A(-1,9), B(3,-3), and L2: C(-1,-1), D(2,5) Are the numbers good?

OpenStudy (anonymous):

yes

OpenStudy (mathmate):



OpenStudy (mathmate):

So we know the answer is expected at about (2,3).

OpenStudy (mathmate):

Do you know how to find the slope between two given points?

OpenStudy (anonymous):

substute x y to find the slope

OpenStudy (mathmate):

Yes, but this way: Slope (m) between two points \(R(x_1,y_1)\) and \(S(x_2,y_2)\) \(\Large m=\frac{y_2-y_1}{x_2-x_1}\)

OpenStudy (mathmate):

We have two lines, and we need the intersection of L1 and L2

OpenStudy (anonymous):

would L1 be 4/1 n L2 be 3/1?

OpenStudy (mathmate):

I have m1=(9--3)/(-1-3)=-3 can you work out m2? I don't have 3/1, remember (5--1) /(2--1)=6/3

OpenStudy (mathmate):

@sciencenerdz11 Still there? or gone to bed?

OpenStudy (anonymous):

sorry im thinkin n workin it out on paper

OpenStudy (anonymous):

l2 x=-1/2 y=-1/5

OpenStudy (anonymous):

im ganna have to go and I cant re enter this so plz help me fast

OpenStudy (anonymous):

bye thx for the help btw

OpenStudy (mathmate):

Well, the slopes for the two lines L1 and L2 are -3 and 2 (according to me, as worked out above). Let's call them m1 and m2.

OpenStudy (mathmate):

You will need to do the following steps: 1. Find the lines L1 and L2 in the form y=m1(x)+b1 y=m2(x)+b2 Details to follow. 2. Find the intersection by solving the above system of 2 linear equations. You can solve by substitution, comparison, or by elimination.

OpenStudy (mathmate):

To find the equations of L1 and L2, you use the slope m (m1 or m2) and a point (A or C) and following steps: [sorry I was away for a little while] Point-slope form of equation of a line \(\color{teal}{m} \)= slope \((\color{blue}{x_0},\color{purple}{y_0})\) = given point Equation of line \(\boxed{\color{red}{L: y-\color{purple}{y_0} = \color{teal}{m} (x-\color{blue}{x_0})}}\) Example: Find equation of the line with slope 2.5 passing through the point (4,3) \(\color{teal}{m} = \color{teal}{2.5}\) \((\color{blue}{x_0},\color{purple}{y_0})=(\color{blue}{4},\color{purple}{3})\) \(L: (y-\color{purple}{3}) = \color{teal}{2.5} (x-\color{blue}{4}) => y-\color{purple}{3}=\color{teal}{2.5}x-10 => y = 2.5x-7\)

OpenStudy (mathmate):

A(-1,9) would be the (x0,y0) for L1, and C(-1,-1) would be the (x0,y0) for L2.